OpenStudy (anonymous):

integral of (3x+7)/x^2

OpenStudy (anonymous):

could u show me the steps. i tried using u=x^2 is that right

OpenStudy (anonymous):

no

OpenStudy (anonymous):

easier to divide each term by x^2 and do it piece by piece

OpenStudy (anonymous):

thats hat i thought

OpenStudy (anonymous):

\[\int\frac{3}{x}dx+\int\frac{7}{x^2}dx\] etc

OpenStudy (anonymous):

thanks so much lemme do it and then we could compare answers

OpenStudy (anonymous):

Yes, actually, you can do it via u substitution. Set u=3x+7, x=(u-7)/3. Which is ugly.

OpenStudy (anonymous):

i'll stick to the first suggestion

OpenStudy (anonymous):

this is my first time on this site...its pretty awesome

OpenStudy (anonymous):

should get the answer easily now yes?

OpenStudy (anonymous):

just to check your answer ... you can click the button "Show steps"

OpenStudy (anonymous):

@mathg8 i tried wolfram but their process is way too long

OpenStudy (anonymous):

even without wolfram i bet you can do it in your head. first part gives \[3\ln(x)\] second give \[-\frac{7}{x}\]

OpenStudy (anonymous):

true ...they're using u substitution ...

OpenStudy (anonymous):

satellite 73 method ...best !

OpenStudy (anonymous):

Int (3x+7)/x^2 = 3 Int ( dx/x) + 7 Int ( dx/x^2) = 3lnx - 7/x + C

OpenStudy (anonymous):

yeah thats what i got

OpenStudy (anonymous):

Do you have any other questions?

OpenStudy (anonymous):

im going through my review for my test tomorrow. if i run into any difficulties i'll let u know

OpenStudy (anonymous):

I'm glad you get the right answer :)

OpenStudy (anonymous):

integral of (ln(ln(x))/(xln(x))

OpenStudy (anonymous):

Let u = lnx -> du = dx/x -> int ( Ln u du/ u ) Again let t = Lnu -> dt = du/u -> Int ( tdt) = t^2/2 = (Lnu)^2/2 = Ln (lnx)^2/ 2 + C

OpenStudy (anonymous):

why are you using a "t" variable